# LLM-TabFlow: Synthetic Tabular Data Generation with Inter-column Logical Relationship Preservation

### v1

@@ 5.3 Inter-Column Relationship Preservation
Retailing Hierarchical Consistency Geographical (Orders)
@@
Mathematical Dependency Sales Values $x_{1,j}$ (quantity), $x_{2,j}$ (price), $x_{3,j}$ (discount rate), $x_{4,j}$ (discount value), $x_{5,j}$ (original price), $x_{6,j}$ (sales price) $D_{2,j}:x_{5,j}=x_{1,j}\times x_{2,j}$
@@
Temporal Dependency Delivery Time $x_{3,j}$ (order created date), $x_{4,j}$ (planned delivery time), $x_{5,j}$ (realistic delivery time) $D_{1,j}:x_{3,j}<x_{4,j}$
@@
Mathematical Dependency Purchase Values $x_{6,j}$ (quantity), $x_{7,j}$ (price), $x_{8,j}$ (discount rate), $x_{9,j}$ (net amount), $x_{10,j}$ (gross amount) $D_{3,j}:x_{9,j}=x_{6,j}\times x_{7,j}$
@@
Retailing HCS 98.09$\pm$0.01 39.23$\pm$0.03 16.07$\pm$0.01 71.63$\pm$0.08 98.01$\pm$0.01 100.00$\pm$0.00
MDI 87.03$\pm$0.03 38.87$\pm$0.05 59.08$\pm$0.05 68.34$\pm$0.08 97.37$\pm$0.02 100.00$\pm$0.00
DSI 74.94$\pm$0.00 30.55$\pm$0.01 79.72$\pm$0.02 66.94$\pm$0.01 98.35$\pm$0.01 98.36$\pm$0.01
Purchasing HCS 81.17$\pm$0.29 0.00$\pm$0.00 0.01$\pm$0.03 54.99$\pm$0.23 99.92$\pm$0.01 100.00$\pm$0.00
MDI 84.82$\pm$0.09 36.69$\pm$0.06 63.49$\pm$0.13 56.30$\pm$0.11 94.54$\pm$0.06 100.00$\pm$0.00
DSI 99.75$\pm$0.02 59.39$\pm$0.01 96.62$\pm$0.01 93.15$\pm$0.00 94.45$\pm$0.02 96.83$\pm$0.02
Through LLM reasoning, the inter-column relationships in the two datasets are identified, as presented in Table 3. These relationships are categorized into different types, with corresponding attributes and conditions specified for both datasets. To quantify the inter-column consistency and dependency, we employ three evaluation metrics—HCS, MDI, and DSI—proposed in Long et al. (2025b). The results are presented in Table 4. As shown, LLM-TabFlow achieves a perfect value of 100% in both consistency and dependency, demonstrating its ability to fully preserve the logical relationships within the real data. GReaT performs slightly worse, exhibiting strong inter-column relationship preservation but with some errors in capturing dependencies across multiple columns. In contrast, latent space-based generative models (e.g., CTGAN, TabDDPM, and TabSyn) perform significantly worse, particularly in maintaining hierarchical consistency across both datasets. Interestingly, SMOTE outperforms most advanced generative models in preserving inter-column logical relationships. To further assess dependency preservation, we evaluate the DSI score by analyzing previously unseen relationships between variables and target values. Specifically, we examine financial attributes to assess their impact on target outcomes across both datasets. LLM-TabFlow achieves the overall best performance, while SMOTE and GReaT perform well on individual datasets. These results highlight LLM-TabFlow’s strength in preserving both simple logical relationships and complex inter-column dependencies within tabular data, setting a new benchmark for synthetic tabular data generation with preserved inter-column relationships.
@@ 5.4 Utility: Machine Learning Efficiency
While data utility, as illustrated in Figure 3, consists of two sub-dimensions for evaluation, this experiment mainly focuses on quantitatively assessing the utility of the generated synthetic data through its machine learning efficiency. The experimental results in Table 5 demonstrate that LLM-TabFlow achieves the overall best performance across both datasets, obtaining the highest AUC on the Retailing dataset and the second-highest on the Purchasing dataset. Specifically, in the Retailing dataset, our approach outperforms TabSyn in terms of AUC scores for classification tasks. GReaT is just behind TabSyn but surpasses SMOTE, demonstrating the advantages of LLM-based training and sampling in generating synthetic tabular data. In contrast, GANs and TabDDPM models are considerably less effective in improving classification efficiency with large-scale synthetic data. However, all these methods still lag significantly behind those trained on the real dataset. In the Purchasing dataset, SMOTE achieves highest AUC score and achieves performance close to that of real data. LLM-TabFlow, which achieves the second-best AUC, along with other methods, still lags behind real data, highlighting the ongoing challenge of achieving parity with real data in machine learning efficency, despite the advancements in generative modeling techniques. Moreover, the ability of synthetic data to preserve inter-column relationships directly impacts its utility in real-world applications, such as process simulation and decision-making in finance and delivery of supply chains. Based on the results in Table 4 and Table 5, LLM-TabFlow generates synthetic tabular data that can effectively substitute real data, achieving satisfactory performance in real-world machine learning tasks.
@@ 5.5 Privacy Preservation
We further evaluate the proposed method alongside the five baseline approaches in the privacy dimension. As described in Section 5.5, DCR (see Equation 10)—a metric that quantifies the similarity between synthetic and real data—is used to evaluate privacy preservation. Additionally, since balancing privacy and utility is crucial for effective tabular data generation, we adopt the Classifier Two-Sample Test (C2ST) (see Section 4.3.1) to measure the distinguishability of real and synthetic data. For this experiment, a logistic regression classifier is used to compute the C2ST. The results in Table 6 show that LLM-TabFlow and TabDDPM achieve the best privacy-preserving performance (highest DCR values) on the Purchasing and Retailing datasets, respectively. However, TabDDPM obtains an extremely low C2ST value, indicating that its strong privacy preservation performance comes at the expensive cost of accuracy. In contrast, SMOTE and LLM-TabFlow achieve the highest and second-highest C2ST scores on the Retailing and Purchasing dataset, respectively, demonstrating strong accuracy in data synthesis. While SMOTE shows the best accuracy in terms of accuracy (highest C2ST value), it achieves the worst (lowest DCR value) privacy preservation performance, highlighting its high accuracy cost at preserving privacy. Our method, however, provides the best balance between privacy and accuracy, outperforming both SMOTE and state-of-the-art generative models. This balance is crucial for applications that require both robust data privacy protection and high data utility, making our approach particularly well-suited for real-world industrial scenarios that demand both privacy guarantees and accurate data synthesis.
@@ 6 Discussion and Implications
We propose a novel approach—LLM-TabFlow—that integrates LLM reasoning with score-based diffusion models to generate synthetic tabular data. This approach significantly outperforms traditional diffusion-based models in capturing complex inter-column logical relationships. Unlike conventional methods that embed all features with intricate relationships directly into latent space, LLM-TabFlow introduces an innovative framework that compresses complex tabular data into simplified representations for generation. These representations are subsequently decompressed back into their original form after sampling. This approach not only enhances the model’s ability to handle complex relationships but also provides a more efficient and interpretable framework for tabular data synthesis. Experimental results demonstrate that PLM-based models excel at capturing inter-column relationships, which may due to their autoregressive nature. However, these methods struggle to accurately preserve the low- and high-order statistics of the original (real) dataset and require approximately ten times more training and inference time compared to latent space-based models. Additionally, it was observed that GReaT can generate unseen categories for product names, revealing the inherent uncertainty of LLM-based generative models. This finding underscores the need for caution when using such models to generate complex categorical data. In contrast, our approach mitigates the uncertainties, providing a more reliable method for tabular data generation. It outperforms existing latent-space-based models in accurately handling datasets with diverse logical dependencies. Furthermore, LLM-TabFlow enhances computational efficiency and improves the capture of heterogeneous data distributions in latent space compared to other LLM-based models. Overall, LLM-TabFlow achieves a much better balance between accuracy, utility, and privacy, positioning it as a robust solution for synthetic tabular data generation.
@@
However, this work has some limitations. Firstly, the identification of relationships in the data is influenced by the knowledge of the LLMs. There is no guarantee that LLMs can capture all complex relationships, particularly when column names are abbreviated or the descriptions are ambiguous or inaccurate. This reliance on LLMs may introduce biases or variances in understanding the underlying data structure. Secondly, the experimental evaluation is limited to datasets within the supply chain domain. To ensure broader applicability, our method needs to be tested on more diverse and comprehensive tabular datasets from other industrial sectors, such as healthcare, finance, and social sciences. This would help verify the generalizaubility of LLM-TabFlow across different types of domain knowledge and data structures. Additionally, privacy concerns arise if column names and descriptions contain sensitive information, as LLMs process these inputs to infer relationships. This could potentially expose private information, especially in domains such as healthcare or finance.
@@
Despite these limitations, LLM-TabFlow shows significant potential for real-world uses across various industries. Beyond preserving data privacy and enhancing machine learning tasks, it is, to the best of our knowledge, the first framework capable of capturing and replicating complex real-world dependencies in synthetic tabular data. This capability is particularly valuable for conducting simulations in digital twins (Xu et al., 2025), enabling more accurate business analysis, predictive modeling, and data-driven decision-making. For example, LLM-TabFlow can be applied to model complex relational data across multiple competing business entities, addressing the data scarcity problems in industrial domains caused by privacy and reluctance to share data (Xu et al., 2025; AlMahri et al., 2024). In agent-based simulations, LLM-TabFlow would generates synthetic data that accurately reflects the interactions between agents, enabling researchers to study complex systems such as supply chains (Xu et al., 2024b, 2025), financial markets (Proselkov et al., 2024), and social networks. Similarly, high-quality, realistic synthetic data would be used to train and fine-tune autonomous agents for customized scenarios, enhancing their capabilities in dynamic environments, such as managing supply chain operations and making sophisticated decision making in autonomous supply chains (Xu et al., 2024a, b, c). These industrial implications highlight LLM-TabFlow’s significant impact in generating high-fidelity, privacy-preserved synthetic data, positioning it as a transformative tool for driving data-driven innovation in both industry and academia.
@@ 7 Conclusion and Future Work
This paper addresses the challenge of preserving inter-column relationships in synthetic tabular data generation for real-world applications. We propose an evaluation framework comprising six dimensions to evaluate the performance of synthetic tabular data in industrial contexts, where both high utility and strong privacy preservation are essential. We have proposed LLM-TabFlow, a novel synthetic tabular data generation approach that integrates LLM reasoning with latent space-based data generation. Additionally, we conduct extensive experiments to evaluate the performance of the proposed method and the other five baselines under the proposed evaluation framework. These experiments show that existing methods struggle to preserve logical inter-column relationships, often failing to maintain dependencies and consistenc. In contrast, effectively upholds hierarchical consistency and logical dependencies while achieving the best trade-off among data fidelity, privacy, and utility.
@@
To the best of our knowledge, LLM-TabFlow is the first solution capable of synthesizing tabular data while simultaneously preserving inter-column relationships and ensuring privacy. This highlights its potential for generating high-quality data for industrial applications, including simulation and advanced decision-making. Moreover, this approach lays the foundation for integrating LLM reasoning with tabular data generation, offering a pipeline that can be extended to other domains such as finance and healthcare. By doing so, LLM-TabFlow unlocks new opportunities for both research and real-world applications, fostering innovation in data-driven industries.
@@
However, our study has certain limitations, as discussed in Section 6. Addressing these limitations will be a key focus of our future work. First, we will evaluate LLM-TabFlow on a broader range of diverse and challenging datasets, including those from various industrial domains and datasets with abbreviated or ambiguous column names and descriptions. This will help assess the model’s adaptability and its ability to capture complex inter-column relationships across various contexts. Second, we will explore to incorporate structural constraints to achieve more precise and context-aware data synthesis. Another key focus of our future work is evaluating LLM-TabFlow’s ability to identify inter-column relationships when sensitive information in column names or descriptions is anonymized. Furthermore, while LLM-TabFlow effectively preserves logical relationships in complex tabular data, real-world tables often contain domain-specific constraints that are not always apparent, such as environmental limits (e.g., industrial temperature ranges) and natural rules (e.g., age must be over 18 for adulthood). Incorporating new prompting methods into our model is crucial for maintaining these structured constraints, ensuring more precise and context-aware data synthesis. Additionally, preserving column relationships across multiple tables is essential for generating realistic synthetic tabular data, especially in scenarios involving foreign key dependencies. Our future work will focus on identifying and maintaining these cross-table relationships to ensure consistency, which is critical for information systems in industries such as supply chain management, e-commerce, and healthcare.

### v2

@@ 7.3 Inter-Column Relationship Preservation
Retailing
Hierarchical
Consistency
Geographical (Orders)
@@
Mathematical
Dependency
Sales Values
$x_{1,j}$ (quantity), $x_{2,j}$ (price),
$x_{3,j}$ (discount rate), $x_{4,j}$ (discount value),
$x_{5,j}$ (original price), $x_{6,j}$ (sales price)
$D_{2,j}:x_{5,j}=x_{1,j}\times x_{2,j}$
@@
Temporal
Dependency
Delivery Time
$x_{3,j}$ (order created date), $x_{4,j}$ (planned delivery time),
$x_{5,j}$ (realistic delivery time)
$D_{1,j}:x_{3,j}<x_{4,j}$
@@
Mathematical
Dependency
Purchase Values
$x_{6,j}$ (quantity), $x_{7,j}$ (price),
$x_{8,j}$ (discount rate), $x_{9,j}$ (net amount), $x_{10,j}$ (gross amount)
$D_{3,j}:x_{9,j}=x_{6,j}\times x_{7,j}$
@@
[…]
Through LLM reasoning, the inter-column relationships in the two datasets are identified, as presented in Table 3. These relationships are categorized into different types, with corresponding attributes and conditions specified for both datasets. To quantify the inter-column consistency and dependency, we employ three evaluation metrics—HCS, MDI, and DSI—proposed in Long et al. (2025b). The results are presented in Table 4. As shown, LLM-TabLogic achieves a perfect value of 100% in both consistency and dependency, demonstrating its ability to fully preserve the logical relationships within the real data. GReaT performs slightly worse, exhibiting strong inter-column relationship preservation but with some errors in capturing dependencies across multiple columns. In contrast, latent space-based generative models (e.g., CTGAN, TabDDPM, and TabSyn) perform significantly worse, particularly in maintaining hierarchical consistency across both datasets. Interestingly, SMOTE outperforms most advanced generative models in preserving inter-column logical relationships. To further assess dependency preservation, we evaluate the DSI score by analyzing previously unseen relationships between variables and target values. Specifically, we examine financial attributes to assess their impact on target outcomes across both datasets. LLM-TabLogic achieves the overall best performance, while SMOTE and GReaT perform well on individual datasets. These results highlight LLM-TabLogic’s strength in preserving both simple logical relationships and complex inter-column dependencies within tabular data, setting a new benchmark for synthetic tabular data generation with preserved inter-column relationships.
@@ 7.4 Utility: Machine Learning Efficiency
While data utility, as illustrated in Figure 3, consists of two sub-dimensions for evaluation, this experiment mainly focuses on quantitatively assessing the utility of the generated synthetic data through its machine learning efficiency. The experimental results in Table 5 demonstrate that LLM-TabLogic achieves the overall best performance across both datasets, obtaining the highest AUC on the Retailing dataset and the second-highest on the Purchasing dataset. Specifically, in the Retailing dataset, our approach outperforms TabSyn in terms of AUC scores for classification tasks. GReaT is just behind TabSyn but surpasses SMOTE, demonstrating the advantages of LLM-based training and sampling in generating synthetic tabular data. In contrast, GANs and TabDDPM models are considerably less effective in improving classification efficiency with large-scale synthetic data. However, all these methods still lag significantly behind those trained on the real dataset. In the Purchasing dataset, SMOTE achieves highest AUC score and achieves performance close to that of real data. LLM-TabLogic, which achieves the second-best AUC, along with other methods, still lags behind real data, highlighting the ongoing challenge of achieving parity with real data in machine learning efficency, despite the advancements in generative modeling techniques. Moreover, the ability of synthetic data to preserve inter-column relationships directly impacts its utility in real-world applications, such as process simulation and decision-making in finance and delivery of supply chains. Based on the results in Table 4 and Table 5, LLM-TabLogic generates synthetic tabular data that can effectively substitute real data, achieving satisfactory performance in real-world machine learning tasks.
@@ 7.5 Privacy Preservation
We further evaluate the proposed method alongside the five baseline approaches in the privacy dimension. As described in Section 7.5, DCR (see Equation 14)—a metric that quantifies the similarity between synthetic and real data—is used to evaluate privacy preservation. Additionally, since balancing privacy and utility is crucial for effective tabular data generation, we adopt the Classifier Two-Sample Test (C2ST) (see Section 5.3.1) to measure the distinguishability of real and synthetic data. For this experiment, a logistic regression classifier is used to compute the C2ST. The results in Table 6 show that LLM-TabLogic and TabDDPM achieve the best privacy-preserving performance (highest DCR values) on the Purchasing and Retailing datasets, respectively. However, TabDDPM obtains an extremely low C2ST value, indicating that its strong privacy preservation performance comes at the expensive cost of accuracy. In contrast, SMOTE and LLM-TabLogic achieve the highest and second-highest C2ST scores on the Retailing and Purchasing dataset, respectively, demonstrating strong accuracy in data synthesis. While SMOTE shows the best accuracy in terms of accuracy (highest C2ST value), it achieves the worst (lowest DCR value) privacy preservation performance, highlighting its high accuracy cost at preserving privacy. Our method, however, provides the best balance between privacy and accuracy, outperforming both SMOTE and state-of-the-art generative models. This balance is crucial for applications that require both robust data privacy protection and high data utility, making our approach particularly well-suited for real-world industrial scenarios that demand both privacy guarantees and accurate data synthesis.
@@ 8 Discussion and Implications
We propose a novel approach—LLM-TabLogic—that integrates LLM reasoning with score-based diffusion models to generate synthetic tabular data. This approach significantly outperforms traditional diffusion-based models in capturing complex inter-column logical relationships. Unlike conventional methods that embed all features with intricate relationships directly into latent space, LLM-TabLogic introduces an innovative framework that compresses complex tabular data into simplified representations for generation. These representations are subsequently decompressed back into their original form after sampling. This approach not only enhances the model’s ability to handle complex relationships but also provides a more efficient and interpretable framework for tabular data synthesis. Experimental results demonstrate that PLM-based models excel at capturing inter-column relationships, which may due to their autoregressive nature. However, these methods struggle to accurately preserve the low- and high-order statistics of the original (real) dataset and require approximately ten times more training and inference time compared to latent space-based models. Additionally, it was observed that GReaT can generate unseen categories for product names, revealing the inherent uncertainty of LLM-based generative models. This finding underscores the need for caution when using such models to generate complex categorical data. In contrast, our approach mitigates the uncertainties, providing a more reliable method for tabular data generation. It outperforms existing latent-space-based models in accurately handling datasets with diverse logical dependencies. Furthermore, LLM-TabLogic enhances computational efficiency and improves the capture of heterogeneous data distributions in latent space compared to other LLM-based models. Overall, LLM-TabLogic achieves a much better balance between accuracy, utility, and privacy, positioning it as a robust solution for synthetic tabular data generation.
@@
However, this work has some limitations. Firstly, the identification of relationships in the data is influenced by the knowledge of the LLMs. There is no guarantee that LLMs can capture all complex relationships, particularly when column names are abbreviated or the descriptions are ambiguous or inaccurate. This reliance on LLMs may introduce biases or variances in understanding the underlying data structure. Secondly, the experimental evaluation is limited to datasets within the supply chain domain. To ensure broader applicability, our method needs to be tested on more diverse and comprehensive tabular datasets from other industrial sectors, such as healthcare, finance, and social sciences. This would help verify the generalizaubility of LLM-TabLogic across different types of domain knowledge and data structures. Additionally, privacy concerns arise if column names and descriptions contain sensitive information, as LLMs process these inputs to infer relationships. This could potentially expose private information, especially in domains such as healthcare or finance.
@@
Despite these limitations, LLM-TabLogic shows significant potential for real-world uses across various industries. Beyond preserving data privacy and enhancing machine learning tasks, it is, to the best of our knowledge, the first framework capable of capturing and replicating complex real-world dependencies in synthetic tabular data. This capability is particularly valuable for conducting simulations in digital twins (Xu et al., 2025), enabling more accurate business analysis, predictive modeling, and data-driven decision-making. For example, LLM-TabLogic can be applied to model complex relational data across multiple competing business entities, addressing the data scarcity problems in industrial domains caused by privacy and reluctance to share data (Xu et al., 2025; AlMahri et al., 2024). In agent-based simulations, LLM-TabLogic would generates synthetic data that accurately reflects the interactions between agents, enabling researchers to study complex systems such as supply chains (Xu et al., 2024b, 2025), financial markets (Proselkov et al., 2024), and social networks. Similarly, high-quality, realistic synthetic data would be used to train and fine-tune autonomous agents for customized scenarios, enhancing their capabilities in dynamic environments, such as managing supply chain operations and making sophisticated decision making in autonomous supply chains (Xu et al., 2024a, b, c). These industrial implications highlight LLM-TabLogic’s significant impact in generating high-fidelity, privacy-preserved synthetic data, positioning it as a transformative tool for driving data-driven innovation in both industry and academia.
@@ 9 Conclusion and Future Work
This paper addresses the challenge of preserving inter-column relationships in synthetic tabular data generation for real-world applications. We propose an evaluation framework comprising six dimensions to evaluate the performance of synthetic tabular data in industrial contexts, where both high utility and strong privacy preservation are essential. We have proposed LLM-TabLogic, a novel synthetic tabular data generation approach that integrates LLM reasoning with latent space-based data generation. Additionally, we conduct extensive experiments to evaluate the performance of the proposed method and the other five baselines under the proposed evaluation framework. These experiments show that existing methods struggle to preserve logical inter-column relationships, often failing to maintain dependencies and consistenc. In contrast, effectively upholds hierarchical consistency and logical dependencies while achieving the best trade-off among data fidelity, privacy, and utility.
@@
To the best of our knowledge, LLM-TabLogic is the first solution capable of synthesizing tabular data while simultaneously preserving inter-column relationships and ensuring privacy. This highlights its potential for generating high-quality data for industrial applications, including simulation and advanced decision-making. Moreover, this approach lays the foundation for integrating LLM reasoning with tabular data generation, offering a pipeline that can be extended to other domains such as finance and healthcare. By doing so, LLM-TabLogic unlocks new opportunities for both research and real-world applications, fostering innovation in data-driven industries.
@@
However, our study has certain limitations, as discussed in Section 8. Addressing these limitations will be a key focus of our future work. First, we will evaluate LLM-TabLogic on a broader range of diverse and challenging datasets, including those from various industrial domains and datasets with abbreviated or ambiguous column names and descriptions. This will help assess the model’s adaptability and its ability to capture complex inter-column relationships across various contexts. Second, we will explore to incorporate structural constraints to achieve more precise and context-aware data synthesis. Another key focus of our future work is evaluating LLM-TabLogic’s ability to identify inter-column relationships when sensitive information in column names or descriptions is anonymized. Furthermore, while LLM-TabLogic effectively preserves logical relationships in complex tabular data, real-world tables often contain domain-specific constraints that are not always apparent, such as environmental limits (e.g., industrial temperature ranges) and natural rules (e.g., age must be over 18 for adulthood). Incorporating new prompting methods into our model is crucial for maintaining these structured constraints, ensuring more precise and context-aware data synthesis. Additionally, preserving column relationships across multiple tables is essential for generating realistic synthetic tabular data, especially in scenarios involving foreign key dependencies. Our future work will focus on identifying and maintaining these cross-table relationships to ensure consistency, which is critical for information systems in industries such as supply chain management, e-commerce, and healthcare.
10 Data Availability Statement
The datasets used in this study include both public and restricted-access sources. The UCI Adult Income Dataset (Becker and Kohavi, 1996) and the DataCo Supply Chain Dataset (Wan, 2021) are publicly available. The MIMIC-III Clinical Database (Johnson et al., 2016b) requires credentialed access via PhysioNet due to privacy constraints. The Purchasing Dataset is proprietary and cannot be shared. Public datasets are cited in the references.